Classic Golf Inc. manages five courses in the Jacksonville, Florida, area. The director of golf wishes to study the number of rounds of golf played per weekday at the five courses. He gathered the following sample information. You will need to calculate the average for the five days of the week. This average is the "expected frequency". You are comparing observed frequencies versus the expected frequency. Day Rounds Monday 125 Tuesday 140 Wednesday 120 Thursday 120 Friday 120 At the 0.01 significance level, is there a difference in the number of rounds played by day of the week? H0: Rounds played is the same for each day. H1: Rounds played is not the same. State the decision rule, using the 0.01 significance level. (Round your answer to 3 decimal places.) Compute the value of chi-square. (Round your answer to 3 decimal places.) What is your decision regarding H0?
Answer #1

\alpha = 0.01

ndf = n - 1 = 5 - 1=4

From Table, critical value of \chi ^{2}= 13.2767

Rejection Rule:
Reject H0 if \chi ^{2} > 13.2767.

Test statistic \chi ^{2} is got as follows:

Day Observed (O) Expected (E) (O - E)2/E
Monday 125 625/5=125 0
Tuesday 140 625/5=125 1.800
Wednesday 120 625/5=125 0.200
Thursday 120 625/5=125 0.200
Friday 120 625/5=125 0.200
Total = \chi ^{2} = 2.400

Since calculated value of \chi ^{2} = 2.400 is less than critical value of \chi ^{2} = 13.2767, the difference is not significant. Fail to reject null hypothesis.

Conclusion:

There is no significant difference in the number of rounds played by day of the week.